OpenStudy (anonymous):

Write the slope-intercept equation for the line with slope -3/5 and y-intercept (0, -1) and graph the line.

OpenStudy (saifoo.khan):

5y = 3x -5

OpenStudy (anonymous):

show all work plz

OpenStudy (saifoo.khan):

y = mx +c y = 3/5x -1 (take lcm) y = (3x - 5)/5 5y = 3x -5